shift summary for day shift 6.13.99

Jim Graham (jagraham@hep.uchicago.edu)
Sun, 13 Jun 1999 17:30:45 -0500

Shift Summary for day shift, 06.13.99

On Shift: JAG, BC

Hours of beam delivered at all intensities: 3.5 hrs
Hours of beam delivered at nominal intensity: 0.0
Hours of E832 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 3.5

Accounting of lost time: no beam

========================================================================
Runs ended this shift:
Run 13557 - BEAM_832, beam studies.
Run 13558 - BEAM_832, beam studies.

Tapes brought to FCC this shift: QKC001-003, QKC007-009,
QKS001-003, QKS007-009, QKK001-003, QKB001, QKB003,
QKLA01-02, QKCR01, QKMU01-04.

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Beam studies by Rick^2.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13557 inherit run with no beam from Rick Kessler and
Ruth Pordes. They have started this run with
the accidental file writing to disk, which is
postulated to cause the DA to hang. The run carried
on for ~5 hrs without hanging.

Ruth wants us to run >reaper_dbg from the da_ktev
window if the run hangs.

80 8:53 Beam arrives. The spill structure is bad, (BOA ~4 via
the acc90-v0 plot)
9:27 beam disappears
11:07 beam returns
11:30 onsplit asks us to split out QKMU04
11:40 Rick Coleman begins beam studies for the target
position and the collimator.

12:34 start looking at online histograms
- DC: sods look ~50-100um high in DC1,3,4
- TRG: big ped swings on MAX(MA) plot
- HA: mean ADC profile out of range.
13557 13:27 stop run to reset for Rick Coleman
13558 13:50 start run

14:45 DC 1Y trips for no apparent reason (like a big
voltage spike). Page Patti and have her run
through the proceedure. This proceedure needs to
be posted near the droeges. Also, the audio alarm
was not at all noticable, and Marj's audio alarm
does not play in the control room. This needs to
be fixed.

15:10 Summary on Beam Studies -- Rick Coleman
1. Verified that target seed (NM2SEED2) is low by 1mm wrt target. Used
NM1U to sweep beam across target and plotted the 90deg monitor and
reg rates normalized to NM2SEM
2. Raised the primary beam by 1mm (8 wires) on the target. New target
posted at epicure
3. Found that the primary collimator was clipping the lower portion of
the secondary beam with the old tune.
4. Verified coll. position by moving it.
5. Verified beam was pointing at center of CsI, to within 1mm.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Regards, Jim